### v1

# Equivalence of Families of Polycyclic Codes over Finite Fields

Abstract
We study the equivalence of families of polycyclic codes associated with polynomials of the form $x^{n}-a_{n-1}x^{n-1}-\ldots-a_{1}x-a_{0}$ over a finite field. We begin with the specific case of polycyclic codes associated with a trinomial $x^{n}-a_{\ell}x^{\ell}-a_{0}$ (for some $0<\ell<n$), which we refer to as $\ell$-trinomial codes, after which we generalize our results to general polycyclic codes. We introduce an equivalence relation called $n$-equivalence, which extends the known notion of $n$-equivalence for constacyclic codes [4]. We compute the number of $n$-equivalence classes for this relation and provide conditions under which two families of polycyclic (or $\ell$-trinomial) codes are equivalent. In particular, we prove that when $\gcd(n,n-\ell)=1$, any $\ell$-trinomial code family is equivalent to a trinomial code family associated with the polynomial $x^{n}-x^{\ell}-1$. Finally, we focus on $p^{\ell}$-trinomial codes of length $p^{\ell+r}$, where $p$ is the characteristic of $\mathbb{F}_{q}$ and $r$ an integer, and provide some examples as an application of the theory developed in this paper.
keywords:
Polycyclic codes, trinomial codes, cyclic codes, constacyclic codes, code equivalence, irreducible polynomials, finite fields.
1 Introduction
Coding theory plays a fundamental role in various applications, such as error detection and correction, data transmission, data storage and reliable communication. It involves the study of efficient encoding and decoding methods for transmitting data reliably over noisy channels. Cyclic codes are one of the most important families of linear codes for both theoretical and practical reasons. They establish a key link between coding theory and algebra and their structure often makes them convenient for implementations. Furthermore, many of the known codes with good parameters are cyclic or related to cyclic codes. Cyclic codes were introduced in the 1950s by Prange in [14] as linear codes with the property that the cyclic shift of any codeword is another codeword. They were later on generalized to constacyclic codes in [3] and to polycyclic codes (also known as pseudo-cyclic codes [13]). Like cyclic (and constacyclic) codes, each polycyclic code over a finite field $\mathbb{F}_{q}$ can be described by an ideal of the polynomial ring $\mathbb{F}_{q}[x]/\langle f(x)\rangle,$ where $f(x)$ is a nonzero polynomial in $\mathbb{F}_{q}[x]$. Polycyclic codes constitute constacyclic codes when $f=x^{n}-\lambda,$ for some non-zero $\lambda$ in $\mathbb{F}_{q},$ and its derivatives cyclic codes ($\lambda=1$) and negacyclic codes ($\lambda=-1$). They have received some attention in the literature, see, for example, [2, 7, 11, 18, 19].
In [4] Chen et al. introduced an equivalence relation "$\sim_{n}$" called $n$-equivalence for the nonzero elements of $\mathbb{F}_{q}$ to classify the families of constacyclic codes of length $n$ over $\mathbb{F}_{q}$ represented by the respective polynomial quotient rings: For $\lambda,\mu\in\mathbb{F}_{q}^{*}$, $\lambda\sim_{n}\mu$ means that there exists a nonzero scalar $a\in\mathbb{F}_{q}^{*}$ such that the map $\Phi_{a}$ from the ring $\mathbb{F}_{q}[x]/\langle x^{n}-\mu\rangle$ to $\mathbb{F}_{q}[x]/\langle x^{n}-\lambda\rangle$, defined by $\Phi_{a}(f(x))=f(ax)$ is an $\mathbb{F}_{q}$-algebra isomorphism which is an isometry with respect to the Hamming distance. [^1] This directly implies that each code in $\mathbb{F}_{q}[x]/\langle x^{n}-\mu\rangle$ has an equivalent code in $\mathbb{F}_{q}[x]/\langle x^{n}-\lambda\rangle$ and vice versa. Equivalently, $\lambda\sim_{n}\mu$ if the polynomial $\lambda x^{n}-\mu$ has at least one root in $\mathbb{F}_{q}[x]$. It is easy to relate the generator polynomial of a $\lambda$-constacyclic code $C=\langle g(x)\rangle$ with that of $\Phi_{a}(C)=\langle g(ax)\rangle$. In a recent work, these notions were generalized to the case of skew constacyclic codes over $\mathbb{F}_{q}$ [12], and to the case of constacyclic codes over finite chain rings [6]. Regarding polycyclic codes, in [2] Nuh Aydin et al. studied several properties of trinomial codes and presented several conjectures related to the equivalence and duality of this class of codes. In [19], all the conjectures proposed in [2] were addressed and methods were provided to construct isodual and self-dual polycyclic codes.
In this paper, we continue the study of polycyclic codes by extending the notion of $n$-equivalence to the case of polycyclic codes associated with a polynomial $x^{n}-a_{n-1}x^{n-1}-\ldots-a_{1}x-a_{0}$ over the finite field $\mathbb{F}_{q}$. We begin with the specific case of polycyclic codes associated with the trinomial $x^{n}-a_{\ell}x^{\ell}-a_{0}$, which we refer to as $\ell$-trinomial codes, after which we generalize our results to general polycyclic codes. We compute the number of $n$-equivalence classes and provide conditions under which two polycyclic (or $\ell$-trinomial) code ambient spaces are equivalent. In particular, we prove that when $\gcd(n,n-\ell)=1$, any $\ell$-trinomial code gives rise to an equivalent ambient space associated with the trinomial $x^{n}-x^{\ell}-1$.
The remainder of this paper is organized as follows. Section 2 provides a review of the basic background on polycyclic codes, and we prove some necessary results on binomial polynomials that will be used in our study. In Sections 3, we study the properties of the $(n,\ell)$-equivalence relation and provide conditions under which two $\ell$-trinomial code families are equivalent. In Section 5, we focus on $p^{\ell}$-trinomial codes of length $p^{\ell+r}$, where $p$ is the characteristic of $\mathbb{F}_{q}$ and $r$ an integer. However in section 6, we generalize our results to general polycyclic codes and provide conditions to obtain equivalence between two two families of polycyclic codes. Finally, we provide some examples as an application of the theory developed in this paper.
2 Preliminaries
In this section, we recall some basic definitions and properties of polycyclic codes. Let $\mathbb{F}_{q}$ be the finite field of order $q$ where $q=p^{s}$ for a prime $p$ and a positive integer $s$. A linear code $C$ of length $n$ over $\mathbb{F}_{q}$ is an $\mathbb{F}_{q}$-subspace of $\mathbb{F}_{q}^{n}$. We define the Hamming weight $w_{H}(c)$ as the number of nonzero components of $c=(c_{0},c_{1},\ldots,c_{n-1})\in\mathbb{F}_{q}^{n}$. The Hamming distance $d(c,c^{\prime})$ between two vectors $c$ and $c^{\prime}$ is defined as $d(c,c^{\prime})=\left|\left\{i\mid c_{i}\neq c^{\prime}_{i}\right\}\right|=w_{H}(c-c^{\prime})$. The (minimum) Hamming distance of a code $C$ is defined as
$d(C):=\min\left\{d(c,c^{\prime})\mid c\neq c^{\prime}\right\}.$
It is well known and easy to see that for a linear code $C$ we have $d(C)=w_{H}(C):=\min\left\{w_{{}_{H}}(c)\mid c\in C,c\neq 0\right\}$. By $[n,k,d]_{q}$ we denote a linear code $C$ over $\mathbb{F}_{q}$ of length $n$, dimension $k$, and minimum distance (at least) $d$.
Definition 2.1 (Polycyclic codes).
Let $C$ be a linear code of length $n$ over $\mathbb{F}_{q}$ and $\vec{a}=\left(a_{{}_{0}},a_{{}_{1}},\ldots,a_{{}_{n-1}}\right)\in\mathbb{F}_{q}^{n}$. We say that $C$ is
  1. 1.
    a right polycyclic code with associated vector $\vec{a}$ if for each codeword $\ c=\left(c_{{}_{0}},c_{{}_{1}},\ldots,c_{{}_{n-1}}\right)\in C$ we have $\left(0,c_{{}_{0}},\ldots,c_{{}_{n-2}}\right)+c_{{}_{n-1}}\vec{a}\in C$,
  2. 2.
    a left polycyclic code with associated vector $\vec{a}$ if for each codeword $\ c=\left(c_{{}_{0}},c_{{}_{1}},\ldots,c_{{}_{n-1}}\right)\in C$ we have $\left(c_{{}_{1}},\ldots,c_{{}_{n-1}},0\right)+c_{{}_{0}}\vec{a}\in C$,
  3. 3.
    a bi-polycyclic code with associated vector $\vec{a}$ if it is both a right and a left polycyclic code with associated vector $\vec{a}.$
Remark 2.2.
For any $\lambda\in\mathbb{F}_{q}^{*},$ the $\lambda$-constacyclic codes are the right polycyclic codes associated with the vector $\vec{a}=\left(\lambda,0,\ldots,0\right)$, and the left polycyclic codes associated with the vector $\vec{b}=\left(0,0,\ldots,\lambda\right).$ In particular, cyclic codes ($\lambda=1$) and negacyclic codes ($\lambda=-1$) are special cases of polycyclic codes.
Definition 2.3 (Trinomial codes, [2]).
Let $\vec{a}=\left(a_{{}_{0}},0,\ldots,0,a_{\ell},0,\ldots,0\right)\in\mathbb{F}_{q}^{n},$ with $a_{0}\neq 0$ and $a_{\ell}\neq 0$, for some $0<\ell<n.$ We say that $C$ is a $\ell$-trinomial code of length $n$ over $\mathbb{F}_{q}$ if it is a (right) polycyclic code with associated vector $\vec{a}.$
In this work, we mainly work with right polycyclic codes, which we simply refer to as polycyclic codes. Under the usual identification of vectors with polynomials, each polycyclic code $C$ of length $n$ associated with a vector $\vec{a}$ is seen as an ideal in the polynomial ring $\mathbb{F}_{q}[x]/\langle f(x)\rangle,$ where $f(x)=x^{n}-\vec{a}(x)=x^{n}-a_{n-1}x^{n-1}-\ldots-a_{1}x-a_{0}.$
In the following proposition we collect some basic results on (right) polycyclic codes.
Proposition 2.4 ([10]).
Let $C\subseteq\mathbb{F}_{q}^{n}$ be a polycyclic code with associated vector $\vec{a}=\left(a_{{}_{0}},a_{{}_{1}},\ldots,a_{{}_{n-1}}\right)$. Then we have the following assertions:
  1. 1.
    The set $\varphi(C)$ is an ideal of the polynomial ring $\mathbb{F}_{q}[x]/\langle x^{n}-\vec{a}(x)\rangle,$ with $\vec{a}(x)=a_{n-1}x^{n-1}+\ldots+a_{1}x+a_{0}.$
  2. 2.
    There is a monic polynomial $g(x)\in\mathbb{F}_{q}[x]$ of least degree which divides $x^{n}-\vec{a}(x)$ and $\varphi(C)=\langle g(x)\rangle.$
  3. 3.
    The set $\{g(x),xg(x),\ldots,x^{{n-\deg(g)-1}}g(x)\}$ forms a basis of $\varphi(C)$ and the dimension of $C$ is $n-\deg(g(x)).$
  4. 4.
    A generator matrix $G$ of $C$ is given by :
    $G=\small{\left(\begin{array}[]{c}\varphi^{-1}(g(x))\\ \\ \varphi^{-1}\left(xg(x)\right)\\ \vdots\\ \vdots\\ \varphi^{-1}\left(x^{{k-1}}g(x)\right)\end{array}\right)=\left(\begin{array}[]{cccccccc}g_{{}_{0}}&g_{{}_{1}}&\ldots&g_{{}_{n-k}}&0&\ldots&\ldots&0\\ \\ 0&g_{{}_{0}}&g_{{}_{1}}&\ldots&g_{{}_{n-k}}&0&\ldots&0\\ \vdots&\ddots&\ddots&\ddots&&\ddots&&\vdots\\ \vdots&&\ddots&\ddots&\ddots&&\ddots&\vdots\\ 0&\ldots&&0&g_{{}_{0}}&g_{{}_{1}}&\ldots&g_{{}_{n-k}}\\ \end{array}\right)}$
    where $k=n-\deg(g(x))$ and $g(x)=\displaystyle{\sum_{i=0}^{n-k}}g_{{}_{i}}x^{i}.$
We now give the definition of generator polynomials for polycyclic codes. Like a cyclic code, a right polycyclic code has many generators, but among all its generators there is a special unique one, called the standard generator of $C$. Any other generator of $C$ is a multiple of the standard generator.
Definition 2.5.
Let $C\subseteq\mathbb{F}_{q}^{n}$ be a non-zero (right) polycyclic code of length $n$ over $\mathbb{F}_{q}$. Then the standard generator of $C$ is the monic polynomial $g(x)$ of least degree in $\mathbb{F}_{q}[x]/\langle f(x)\rangle$ such that $\varphi(C)=\langle g(x)\rangle$. In this paper, we refer to the standard generator of $C$ as ”the generator”.
Furthermore, we recall the following lemmas with respect to the number of solutions of a binomial equation in $\mathbb{F}_{q}$ and the degree of $\gcd\left(x^{n}-a,x^{m}-b\right)$.
Lemma 2.6.
[17, Lemma 1] Consider the finite field $\mathbb{F}_{q}$ and $a\in\mathbb{F}_{q}^{*}$. Let $n>0$ be an integer and $d=\gcd(n,q-1)$. Then the equation
$x^{n}-a=0$
has solutions in $\mathbb{F}_{q}$ if and only if
$a^{\frac{q-1}{d}}=1,$
in which case there are exactly $d$ (different) solutions in $\mathbb{F}_{q}.$
Lemma 2.7.
[Redei1950, Lemma 1] Let $m,n\geq 1$ be integers and $a,b\in\mathbb{F}_{q}^{*}$. Then $\gcd\left(x^{n}-a,x^{m}-b\right)$ has degree 0 or $d:=\gcd(m,n)$ (over an arbitrary field). Moreover, $\gcd\left(x^{n}-a,x^{m}-b\right)$ has degree $d$ if and only if $a^{m/d}=b^{n/d}$.
In [1, Theorem 4.1], the authors proved that when $\gcd(n,p)=\gcd(m,p)=1$, the polynomial $\gcd(x^{n}-a,x^{m}-b)$ is either $1$ or $x^{\gcd(n,m)}-c$, for some $c\in\mathbb{F}_{q}$. Following their proof, we extend the result to the general case in the following lemma, allowing also $\gcd(n,p)\neq 1$ or $\gcd(m,p)\neq 1.$
Lemma 2.8.
Let $f=x^{n}-a$ and $g=x^{m}-b$ be two polynomials in $\mathbb{F}_{q}[x]$. Then either $\gcd(f,g)=1$ or
$\gcd(f,g)=x^{\gcd(m,n)}-a^{u}b^{v},$
for some integers $u$ and $v$ such that $\gcd(m,n)=un+vm$.
Proof.
Let $n=p^{r}n^{\prime}$ and $m=p^{s}m^{\prime}$ such that $\gcd(n^{\prime},p)=\gcd(m^{\prime},p)=1$. Then there exists a unique $a^{\prime}=a^{p^{-r}}\in\mathbb{F}_{q}^{*}$ and $b^{\prime}=b^{p^{-s}}\in\mathbb{F}_{q}^{*}$ such that
$f=x^{n}-a=\left(x^{n^{\prime}}-a^{\prime}\right)^{p^{r}},\quad\text{and}\ \ g=x^{m}-b=\left(x^{m^{\prime}}-b^{\prime}\right)^{p^{s}}.$
Let $t=\operatorname{ord}_{n^{\prime}r_{1}m^{\prime}r_{2}}(q)$ be the order of $q$ modulo $n^{\prime}r_{1}m^{\prime}r_{2},$ where $r_{1}$ and $r_{2}$ are the multiplicative orders of $a$ and $b$ in $\mathbb{F}_{q}^{*}$, respectively. Then the field extension $\mathbb{F}_{q^{t}}$ contains both the roots of $x^{n^{\prime}}-a^{\prime}$ and $x^{m^{\prime}}-b^{\prime}$.
Let $\zeta$ be a primitive $(n^{\prime}m^{\prime})^{\text{th}}$ root of unity. Then $\zeta^{n^{\prime}}$ and $\zeta^{m^{\prime}}$ are respectively the $m^{\prime\text{th}}$ and $n^{\prime\text{th}}$ roots of unity. If $\gcd(x^{n^{\prime}}-a^{\prime},x^{m^{\prime}}-b^{\prime})=1,$ then $\gcd(f,g)=1$, and we are done.
Now, suppose there exists a common root $\delta$ of $x^{n^{\prime}}-a^{\prime}$ and $x^{m^{\prime}}-b^{\prime}$, which is an $n^{\prime\text{th}}$ root of $a^{\prime}$ and an $m^{\prime\text{th}}$ root of $b^{\prime}$. As in the proof of [1, Theorem 4.1], the roots of $x^{n^{\prime}}-a^{\prime}$ and $x^{m^{\prime}}-b^{\prime}$ are, respectively,
$\delta,\delta\zeta^{m^{\prime}},\delta(\zeta^{m^{\prime}})^{2},\ldots,\delta(\zeta^{m^{\prime}})^{n^{\prime}-1},\ \ \text{and}\ \ \delta,\delta\zeta^{n^{\prime}},\delta(\zeta^{n^{\prime}})^{2},\ldots,\delta(\zeta^{n^{\prime}})^{m^{\prime}-1}.$
Let $d^{\prime}:=\gcd(n^{\prime},m^{\prime})$ and $\beta:=\zeta^{\operatorname{lcm}{(n^{\prime},m^{\prime})}},$ then the roots of $\gcd(x^{n^{\prime}}-a^{\prime},x^{m^{\prime}}-b^{\prime})$ are
$\delta,\delta\beta,\delta\beta^{2},\ldots,\delta\beta^{d^{\prime}-1},$
and so
$\gcd(x^{n^{\prime}}-a^{\prime},x^{m^{\prime}}-b^{\prime})=x^{d^{\prime}}-\delta^{d^{\prime}}.$
Note that the roots of $f$ are:
$\overbrace{\delta,\ldots,\delta}^{p^{r}\text{ times}},\overbrace{\delta\zeta^{m^{\prime}},\ldots,\delta\zeta^{m^{\prime}}}^{p^{r}\text{ times}},\ldots,\overbrace{\delta(\zeta^{m^{\prime}})^{n^{\prime}-1},\ldots,\delta(\zeta^{m^{\prime}})^{n^{\prime}-1}}^{p^{r}\text{ times}},$
and the roots of $g$ are
$\overbrace{\delta,\ldots,\delta}^{p^{s}\text{ times}},\overbrace{\delta\zeta^{n^{\prime}},\ldots,\delta\zeta^{n^{\prime}}}^{p^{s}\text{ times}},\ldots,\overbrace{\delta(\zeta^{n^{\prime}})^{m^{\prime}-1},\ldots,\delta(\zeta^{n^{\prime}})^{m^{\prime}-1}}^{p^{s}\text{ times}}.$
It follows that the roots of $\gcd(f,g)$ are
$\overbrace{\delta,\ldots,\delta}^{p^{\min(r,s)}\text{ times}},\overbrace{\delta\beta,\ldots,\delta\beta}^{p^{\min(r,s)}\text{ times}},\ldots,\overbrace{\delta\beta^{d^{\prime}-1},\ldots,\delta\beta^{d^{\prime}-1}}^{p^{\min(r,s)}\text{ times}}.$
Therefore, $\deg(\gcd(f,g))=p^{\min(r,s)}d^{\prime},$ and
$\gcd(f,g)=\left(x^{d^{\prime}}-\delta^{d^{\prime}}\right)^{p^{\min(r,s)}}=x^{p^{\min(r,s)}d^{\prime}}-\delta^{d^{\prime}p^{\min(r,s)}}=x^{\gcd(m,n)}-\delta^{\gcd(m,n)}.$
To complete the proof we need to show that $\delta^{d}=a^{u}b^{v}\in\mathbb{F}_{q},$ for some integers $u,v$ such that $d=\gcd(m,n)=un+vm$. We know that $\delta^{n}=a$ and $\delta^{m}=b$, which implies
$\delta^{d}=\delta^{un+vm}=a^{u}b^{v}\in\mathbb{F}_{q}.$
Finally, we proved that $\gcd(f,g)=x^{d}-a^{u}b^{v}$ for some integers $u$ and $v$ such that $d=\gcd(m,n)=un+vm$.
More generally, we prove by induction the following result:
Lemma 2.9.
Let $m$ be an integer and let $f_{i}=x^{n_{i}}-a_{i}$ for $i=1,\ldots,m$ be polynomials in $\mathbb{F}_{q}[x]$. Then $\gcd(f_{1}(x),f_{2}(x),\ldots,f_{m}(x))$ is either $1$ or of the form
$x^{d}-\displaystyle{\prod_{i=1}^{m}}a_{i}^{u_{i}},$
where the $u_{i}$ are integers such that $d=\gcd(n_{1},n_{2},\ldots,n_{m})=\displaystyle{\sum_{i=1}^{m}}u_{i}n_{i}$.
Proof.
Let $f_{i}=x^{n_{i}}-a_{i}$ for $i=1,\ldots,m$. We will use induction to prove that $\gcd(f_{1},f_{2},\ldots,f_{m})$ is either $1$ or of the form $x^{d}-\displaystyle{\prod_{i=1}^{m}}a_{i}^{u_{i}}$, where $d=\gcd(n_{1},n_{2},\ldots,n_{m})$.
For $m=2,$ the result is verified by Lemma 2.8. So, let us assume that the result is valid for the $m-1$ polynomials $f_{1},\ldots,f_{m-1}$. That is,
$\gcd(f_{1},f_{2},\ldots,f_{m-1})=\begin{cases}x^{d^{\prime}}-\displaystyle{\prod_{i=1}^{m-1}}a_{i}^{u_{i}},&\text{if }d^{\prime}=\gcd(n_{1},n_{2},\ldots,n_{m-1})=\displaystyle{\sum_{i=1}^{m-1}}u_{i}n_{i}\\ 1,&\text{else}\end{cases}.$
Let us now consider $f_{1},f_{2},\ldots,f_{m}$, and put $g:=\gcd(f_{1},f_{2},\ldots,f_{m-1})$. If $g=1$, then $\gcd(g,f_{m})=1$. Else, $g=x^{d^{\prime}}-\displaystyle{\prod_{i=1}^{m-1}}a_{i}^{u_{i}}$, then by applying Lemma 2.8 to $g$ and $f_{m}=x^{n_{m}}-a_{m}$, we obtain
$\gcd(g,f_{m})=\begin{cases}x^{d}-\displaystyle{\prod_{i=1}^{m}}a_{i}^{u_{i}},&\text{if there is a common root,}\\ 1,&\text{if there is no common root of }g\text{ and }f_{m},\end{cases}$
where $d=\gcd(d^{\prime},n_{m})=\gcd(n_{1},n_{2},\ldots,n_{m})$, and $u_{i}$ are integers such that $d=\displaystyle{\sum_{i=1}^{m}}u_{i}n_{i}$. Then the result holds.
3 Equivalence of Trinomial Code Families
In this section, we study the equivalence between $\ell$-trinomial code families.
Definition 3.1.
Let $a_{0},a_{\ell},b_{0},b_{\ell}$ be nonzero elements of $\mathbb{F}_{q}$, and let $\ell$ be an integer such that $0<\ell<n$. We say that $(a_{0},a_{\ell})$ and $(b_{0},b_{\ell})$ are $(n,\ell)$-equivalent in $\mathbb{F}_{q}^{*}\times\mathbb{F}_{q}^{*}$, denoted by
$(a_{0},a_{\ell})\sim_{(n,\ell)}(b_{0},b_{\ell}),$
if there exists an $\alpha\in\mathbb{F}_{q}^{*}$ such that the following map
$\begin{array}[]{cccc}\varphi_{\alpha}:&\mathbb{F}_{q}[x]/\langle x^{n}-b_{\ell}x^{\ell}-b_{0}\rangle&\longrightarrow&\mathbb{F}_{q}[x]/\langle x^{n}-a_{\ell}x^{\ell}-a_{0}\rangle,\\ &f(x)&\longmapsto&f(\alpha x),\end{array}$ (1)
is an $\mathbb{F}_{q}$-algebra isomorphism. Note that such an isomorphism preserves the Hamming weight, i.e.,
$d_{H}(\varphi_{\alpha}(f(x)),\varphi_{\alpha}(g(x)))=d_{H}(f(x),g(x)),$
for all $f(x),g(x)\in\mathbb{F}_{q}[x]/\langle x^{n}-b_{\ell}x^{\ell}-b_{0}\rangle.$
Remark 3.2.
  1. 1.
    For any integer $0<\ell<n,$ the relation $\sim_{(n,\ell)}$ is an equivalence relation on $\mathbb{F}_{q}^{*}\times\mathbb{F}_{q}^{*}.$
  2. 2.
    Note that the $(n,\ell)$-equivalence relation in the above definition generalizes the $n$-equivalence of constacyclic codes studied in [4, Definition 3.1], which was denoted by $\lambda\sim_{n}\mu$.
In the following theorem we give essential characterizations of the $(n,\ell)$-equivalence between two classes of $\ell$-trinomial codes of length $n$ over $\mathbb{F}_{q}$. In the statement we will use the component-wise product of two length $n$ vectors $x$ and $y$, also known as the Schur product, defined as
$(x_{0},x_{1},\ldots,x_{n-1})\star(y_{0},y_{1},\ldots,y_{n-1}):=(x_{0}y_{0},x_{1}y_{1},\ldots,x_{n-1}y_{n-1}).$
Theorem 3.3.
Let $0<\ell<n$ be an integer, $(a_{0},a_{\ell})$ and $(b_{0},b_{\ell})$ be elements of $\mathbb{F}_{q}^{*}\times\mathbb{F}_{q}^{*}$, and $\xi$ be a primitive element of $\mathbb{F}_{q}$. The following statements are equivalent:
  1. 1.
    $(a_{0},a_{\ell})\sim_{(n,\ell)}(b_{0},b_{\ell}).$
  2. 2.
    The polynomials $a_{i}x^{n-i}-b_{i}\in\mathbb{F}_{q}[x]$, with $i\in\{0,\ell\}$, have a common root in $\mathbb{F}_{q}^{*}.$
  3. 3.
    The polynomial $\gcd(a_{0}x^{n}-b_{0},a_{\ell}x^{n-\ell}-b_{\ell})$ has at least one root in $\mathbb{F}_{q}^{*}.$
  4. 4.
    The polynomial $\gcd(x^{n}-b_{0}a_{0}^{-1},x^{n-\ell}-b_{\ell}a_{\ell}^{-1})$ has at least one root in $\mathbb{F}_{q}^{*}.$
  5. 5.
    There exists $\alpha\in\mathbb{F}_{q}^{*}$ such that $(a_{0},a_{\ell})\star(\alpha^{n},\alpha^{n-\ell})=(b_{0},b_{\ell})$.
  6. 6.
    $(a_{0},a_{\ell})^{-1}\star(b_{0},b_{\ell})\in H,$ where $H$ is the cyclic subgroup of $\mathbb{F}_{q}^{*}\times\mathbb{F}_{q}^{*}$ generated by $(\xi^{n},\xi^{n-\ell}).$
The equivalence between (1) and (6) implies that the number of $(n,\ell)$-equivalence classes is
$N_{(n,\ell)}:=\dfrac{(q-1)^{2}}{\operatorname{lcm}(\frac{q-1}{\gcd(n,q-1)},\frac{q-1}{\gcd(n-\ell,q-1)})}=(q-1)\gcd\left(\frac{q-1}{\gcd(n,q-1)},\frac{q-1}{\gcd(n-\ell,q-1)}\right).$
Proof.
  1. (1) $\Rightarrow$ (2)
    Suppose that $(a_{0},a_{\ell})\sim_{(n,\ell)}(b_{0},b_{\ell})$. Then – by Definition 3.1 – there exists $\alpha\in\mathbb{F}_{q}^{*}$ such that the map
    $\varphi_{\alpha}:\mathbb{F}_{q}[x]/\langle x^{n}-b_{\ell}x^{\ell}-b_{0}\rangle\to\mathbb{F}_{q}[x]/\langle x^{n}-a_{\ell}x^{\ell}-a_{0}\rangle,\quad f(x)\ \mapsto\ f(\alpha x)$
    is an $\mathbb{F}_{q}$-algebra isometry. It follows that
    $\varphi_{\alpha}(x^{i})=\varphi_{\alpha}(x)^{i}=\alpha^{i}x^{i}\ \mod(x^{n}-a_{\ell}x^{\ell}-a_{0}),\ \forall i=0,1,\ldots,n-1.$
    Since $\varphi_{\alpha}$ is an $\mathbb{F}_{q}$-algebra isometry and $\varphi(x^{n}-b_{\ell}x^{\ell}-b_{0})=0\ \mod(x^{n}-a_{\ell}x^{\ell}-a_{0})$, then
    $\varphi_{\alpha}(x^{n})=b_{\ell}\alpha^{\ell}x^{\ell}+b_{0}.$ (2)
    On the other hand,
    $\varphi_{\alpha}(x^{n})=\alpha^{n}x^{n}=\alpha^{n}(a_{\ell}x^{\ell}+a_{0})=\alpha^{n}a_{\ell}x^{\ell}+\alpha^{n}a_{0}.$ (3)
    Comparing term by term, we deduce that $a_{0}\alpha^{n}=b_{0}$ and $a_{\ell}\alpha^{n-\ell}=b_{\ell}$, which means that $\alpha$ is a common root of the polynomials $a_{0}x^{n}-b_{0}$ and $a_{\ell}x^{n-\ell}-b_{\ell}$.
  2. (2) $\Rightarrow$ (3)
    and (3) $\Rightarrow$ (4) are immediate.
  3. (4) $\Rightarrow$ (5)
    Let $\alpha$ be a root of the polynomial $\gcd(x^{n}-b_{0}a_{0}^{-1},x^{n-\ell}-b_{\ell}a_{\ell}^{-1})$. Then $\alpha$ is a common root of the polynomials $a_{0}x^{n}-b_{0}$ and $a_{\ell}x^{n-\ell}-b_{\ell}$. It follows that $a_{i}\alpha^{n-i}=b_{i},\ \text{for any}\ i\in\{0,\ell\}$, i.e.,
    $(b_{0},b_{\ell})=(\alpha^{n}a_{0},\alpha^{n-\ell}a_{\ell})=(\alpha^{n},\alpha^{n-\ell})\star(a_{0},a_{\ell}).$
  4. (5) $\Rightarrow$ (6)
    Suppose that there is $\alpha\in\mathbb{F}_{q}^{*}$ such that $(b_{0},b_{\ell})=(\alpha^{n},\alpha^{n-\ell})\star(a_{0},a_{\ell}).$ Then,
    $(a_{0},a_{\ell})^{-1}\star(b_{0},b_{\ell})=(a_{0}^{-1}b_{0},a_{\ell}^{-1}b_{\ell})=(\alpha^{n},\alpha^{n-\ell})=(\xi^{jn},\xi^{j(n-\ell)})=(\xi^{n},\xi^{(n-\ell)})^{j}.$
    It follows that $(a_{0},a_{\ell})^{-1}\star(b_{0},b_{\ell})$ belongs to the cyclic subgroup $H$ generated by $(\xi^{n},\xi^{(n-\ell)})$ as a subgroup of $\mathbb{F}_{q}^{*}\times\mathbb{F}_{q}^{*}.$
  5. (6) $\Rightarrow$ (1)
    Suppose that $(a_{0},a_{\ell})^{-1}\star(b_{0},b_{\ell})$ is an element of the cyclic subgroup $H$ generated by $(\xi^{n},\xi^{(n-\ell)})$ as a subgroup of $\mathbb{F}_{q}^{*}\times\mathbb{F}_{q}^{*}$. Then there exists an integer $h$ such that
    $(a_{0},a_{\ell})^{-1}\star(b_{0},b_{\ell})=(\xi^{n},\xi^{n-\ell})^{h}=(\xi^{hn},\xi^{h(n-\ell)})$
    For $\beta=\xi^{h},$ we obtain that $a_{i}\beta^{n-i}=b_{i},$ for any $i\in\{0,k\}.$ Now, let consider the map $\tilde{\varphi}_{\beta},$ as follows:
    $\begin{array}[]{cccc}\tilde{\varphi}_{\beta}:&\mathbb{F}_{q}[x]&\longrightarrow&\mathbb{F}_{q}[x]/\langle x^{n}-a_{\ell}x^{\ell}-a_{0}\rangle,\\ &f(x)&\longmapsto&f(\beta x).\end{array}$ (4)
    $\tilde{\varphi}_{\beta}$ is a surjective $\mathbb{F}_{q}$-algebra homomorphism, indeed, for all $10\leq j\leq n-$ 1, $x^{j}=\tilde{\varphi}_{\beta}(\beta^{-j}x^{j})$. Moreover,
    $\begin{array}[]{rl}\tilde{\varphi}_{\beta}(x^{n}-b_{\ell}x^{\ell}-b_{0})&=\beta^{n}x^{n}-\beta^{\ell}b_{\ell}x^{\ell}-b_{0}\\ &=\beta^{n}x^{n}-\beta^{n}a_{\ell}x^{\ell}-\beta^{n}a_{0}\\ &=\beta^{n}(x^{n}-a_{\ell}x^{\ell}-a_{0})\\ &=0\mod(x^{n}-a_{\ell}x^{\ell}-a_{0}).\end{array}$
    So $\langle x^{n}-b_{\ell}x^{\ell}-b_{0}\rangle\subseteq\ker\tilde{\varphi}_{\beta}$. And for all $f(x)\in\ker\tilde{\varphi}_{\beta}$,
    $f(\beta x)=0\leavevmode\nobreak\ (\text{mod }x^{n}-a_{\ell}x^{\ell}-a_{0}),$
    then there exists $g(x)\in\mathbb{F}_{q}[x]$ such that $f(\beta x)=g(x)(x^{n}-a_{\ell}x^{\ell}-a_{0})$, thus
    $\begin{array}[]{rl}f(x)&=g(\beta^{-1}x)(\beta^{-n}x^{n}-\beta^{-\ell}a_{\ell}x^{\ell}-a_{0})\\ &=\beta^{-n}g(\beta^{-1}x)(x^{n}-\beta^{n-\ell}a_{\ell}x^{\ell}-\beta^{n}a_{0})\\ &=\beta^{-n}g(\beta^{-1}x)(x^{n}-b_{\ell}x^{\ell}-b_{0}),\ \text{since $a_{i}\beta^{n-i}=b_{i},\ \forall i\in\{0,\ell\}$}\\ \end{array}$
    So $\ker\tilde{\varphi}_{\beta}\subseteq\langle x^{n}-a_{\ell}x^{\ell}-a_{0}\rangle$ and hence $ker\tilde{\varphi}_{\beta}=\langle x^{n}-b_{\ell}x^{\ell}-b_{0}\rangle$. Therefore by the first isomorphism theorem the map
    $\begin{array}[]{cccc}\varphi_{\beta}:&\mathbb{F}_{q}[x]/\langle x^{n}-b_{\ell}x^{\ell}-b_{0}\rangle&\longrightarrow&\mathbb{F}_{q}[x]/\langle x^{n}-a_{\ell}x^{\ell}-a_{0}\rangle,\\ &f(x)&\longmapsto&f(\beta x).\end{array}$ (5)
    is an $\mathbb{F}_{q}$-algebra isomorphism. As the weights of $f(x)$ and $f(\beta x)$ are the same, the result holds.
By the equivalence between (1) and (6), we deduce that the number of $(n,\ell)$-equivalence classes on $\mathbb{F}_{q}^{*}\times\mathbb{F}_{q}^{*}$ corresponds to the order of the group $\left(\mathbb{F}_{q}^{*}\times\mathbb{F}_{q}^{*}\right)/H,$ which equals $N_{(n,\ell)}=\dfrac{(q-1)^{2}}{\operatorname{lcm}(\frac{q-1}{\gcd(n,q-1)},\frac{q-1}{\gcd(n-\ell,q-1)})}=(q-1)\gcd\left(\frac{q-1}{\gcd(n,q-1)},\frac{q-1}{\gcd(n-\ell,q-1)}\right).$
Using the equivalence between the assertions $(1)$ and $(5)$ of Theorem 3.3 we derive a characterization regarding the equivalence between the class of $\ell$-trinomial codes associated with $x^{n}-a_{\ell}x^{\ell}-a_{0}$ and the class associated with $x^{n}-x^{\ell}-1$ in the following.
For this we first derive a result from Lemmas 2.7 and 2.8, which we will then use in the proof of Corollary 3.5.
Lemma 3.4.
Let $m,n$ be two positive integers and $u,v\in\mathbb{Z}$ be such that $\gcd(m,n)=un+vm$. Let furthermore $f=x^{n}-a$ and $\ g=x^{m}-b$ two polynomials in $\mathbb{F}_{q}[x]$. Then
$\gcd(f,g)=\begin{cases}x^{\gcd(n,m)}-a^{u}b^{v}&\text{ if }a^{\frac{m}{\gcd(m,n)}}=b^{\frac{n}{\gcd(m,n)}}\\ 1&\text{ else }\end{cases}.$
Corollary 3.5.
  1. 1.
    The class of $\ell$-trinomial codes associated with $x^{n}-a_{\ell}x^{\ell}-a_{0}$ is equivalent to the class of $\ell$-trinomial codes associated with $x^{n}-x^{\ell}-1.$
  2. 2.
    There exists $\alpha\in\mathbb{F}_{q}^{*}$ such that $(a_{0},a_{\ell})\star(\alpha^{n},\alpha^{n-\ell})=(1,1)$.
  3. 3.
    There exists $\alpha\in\mathbb{F}_{q}^{*}$ being an $n$-th root of $a_{0}$ such that $a_{0}=\alpha^{\ell}a_{\ell}$.
  4. 4.
    We have
    $a_{\ell}^{\frac{n}{\gcd(n,n-\ell)}}=a_{0}^{\frac{n-\ell}{{\gcd(n,n-\ell)}}}$
    and $x^{\gcd(n,n-\ell)}-a_{0}^{v}a_{\ell}^{u}$ has a root, for $u,v\in\mathbb{Z}$ with $\gcd(n,n-\ell)=nv+u(n-\ell).$
Proof.
  • (1) $\Rightarrow$ (2)
    Follows from the equivalence of assertions (1) and (5) of Theorem 3.3.
  • (2) $\Rightarrow$ (3)
    If $(a_{0},a_{\ell})\star(\alpha^{n},\alpha^{n-\ell})=(1,1)$, then
    $a_{0}\alpha^{n}=1\quad\text{ and }\quad a_{\ell}\alpha^{n-\ell}=1$
    $\iff a_{0}=\alpha^{-n}\quad\text{ and }\quad a_{\ell}\alpha^{-\ell}=a_{0}$
    i.e., $\beta:=\alpha^{-1}$ is an $n$-th root of $a_{0}$ and $a_{0}=\beta^{\ell}a_{\ell}$.
  • (3) $\Rightarrow$ (4)
    Suppose that $\alpha$ is an $n$-th root of $a_{0}$ such that $a_{\ell}^{-1}a_{0}=\alpha^{\ell}$. It follows that
    $\alpha^{n}a_{0}^{-1}=1\ \ \text{ and }\ \ a_{\ell}^{-1}a_{0}=a_{\ell}^{-1}a_{0}\alpha^{n}a_{0}^{-1}=\alpha^{\ell}.$
    Hence
    $\alpha^{n}=a_{0}\ \ \text{ and }\ \ \alpha^{n-\ell}=a_{\ell},$
    which means that $\alpha$ is a common root of $x^{n}-a_{0}$ and $x^{n-\ell}-a_{\ell}$. Moreover, we get
    $a_{\ell}^{\frac{n}{\gcd(n,n-\ell)}}=\alpha^{\frac{(n-\ell)n}{\gcd(n,n-\ell)}}=a_{0}^{\frac{n-\ell}{\gcd(n,n-\ell)}},$
    i.e., we can use Lemma 3.4 to deduce that $\gcd(x^{n}-a_{0},x^{n-\ell}-a_{\ell})=x^{\gcd(n,n-\ell)}-a_{\ell}^{u}a_{0}^{v},$ for $u,v\in\mathbb{Z}$ with $vn+u(n-\ell)=\gcd(n,n-\ell).$ It follows that $\alpha$ is a root of $x^{\gcd(n,n-\ell)}-a_{\ell}^{u}a_{0}^{v}$.
  • (4) $\Rightarrow$ (1)
    Suppose that $a_{\ell}^{\frac{n}{\gcd(n,n-\ell)}}=a_{0}^{\frac{n-\ell}{{\gcd(n,n-\ell)}}}$ and that $\alpha$ is a root of $x^{\gcd(n,n-\ell)}-a_{0}^{v}a_{\ell}^{u}$. By Lemma 3.4 it follows that $x^{\gcd(n,n-\ell)}-a_{0}^{v}a_{\ell}^{u}=\gcd(x^{n}-a_{0},x^{n-\ell}-a_{\ell})$ and hence that $\alpha$ is a common root of $x^{n}-a_{0}$ and $x^{n-\ell}-a_{\ell.}$ The statement now follows from Theorem 3.3.
Once we know that two classes are equivalent, we can determine how many different $\varphi_{\alpha}$ are isometric isomorphisms from one class to the other:
Corollary 3.6.
With the same notation as in Theorem 3.3 the number of isometric $\mathbb{F}_{q}$-algebra isomorphisms $\varphi_{\alpha}$ between $\mathbb{F}_{q}[x]/\langle x^{n}-b_{\ell}x^{\ell}-b_{0}\rangle$ and $\mathbb{F}_{q}[x]/\langle x^{n}-a_{\ell}x^{\ell}-a_{0}\rangle$ is $\gcd(n,n-\ell,q-1)$, if $(a_{0},a_{\ell})\sim_{(n,\ell)}(b_{0},b_{\ell})$ (otherwise, there are none).
Proof.
By the first step in the proof of Theorem 3.3, $(a_{0},a_{\ell})\sim_{(n,\ell)}(b_{0},b_{\ell})$ under $\varphi_{\alpha}$ if and only if $\alpha$ is a root of the polynomial $\gcd(x^{n}-b_{0}a_{0}^{-1},x^{n-\ell}-b_{\ell}a_{\ell}^{-1})$. By Lemma 2.8, we obtain that
$\gcd(x^{n}-b_{0}a_{0}^{-1},x^{n-\ell}-b_{\ell}a_{\ell}^{-1})=x^{\gcd(n,n-\ell)}-(b_{0}a_{0}^{-1})^{u}(b_{\ell}a_{\ell}^{-1})^{v},$
where $d=\gcd(n,n-\ell)=un+v(n-\ell)$ for some integer $u$ and $v$. By Lemma 2.6, the number of roots of $x^{\gcd(n,n-\ell)}-(b_{0}a_{0}^{-1})^{u}(b_{\ell}a_{\ell}^{-1})^{v}$ in $\mathbb{F}_{q}$ is equal to $\gcd(n,n-\ell,q-1)$, hence there are $\gcd(n,n-\ell,q-1)$ many different $\alpha$ such that $\varphi_{\alpha}$ is an isomorphism. (For the last step note that $\alpha\in\mathbb{F}_{q}^{*}$ since 00 is not a root of the polynomials above.)
In the following result, we describe the associated polynomials of possible $\ell$-trinomial codes based on the $(n,\ell)$-equivalence relation defined above.
Theorem 3.7.
Let $n,\ell$ be two integers such that $0<\ell<n$ and let $\xi$ be a primitive element of $\mathbb{F}_{q}$. Set $d:=\gcd\left(\frac{q-1}{\gcd(n,q-1)},\frac{q-1}{\gcd(n-\ell,q-1)}\right)$ and $d_{i}:=\gcd(n-i,q-1)$, for $\ i\in\{0,\ell\}$. Moreover, let $a_{0},a_{\ell}\in\mathbb{F}_{q}^{*}$.
  1. 1.
    If $d=1,$ then the class of $\ell$-trinomial codes associated with $x^{n}-a_{\ell}x^{\ell}-a_{0}$ is equivalent to the class of $\ell$-trinomial codes associated with $x^{n}-\xi^{j}x^{\ell}-\xi^{i},$ for some $i\in\{0,1,\ldots,d_{0}-1\}$ and $j\in\{0,1,\ldots,d_{\ell}-1\}$.
  2. 2.
    If $d\neq 1,$ then the class of $\ell$-trinomial codes associated with $x^{n}-a_{\ell}x^{\ell}-a_{0}$ is equivalent to the class of $\ell$-trinomial codes associated with $x^{n}-\xi^{j}x^{\ell}-\xi^{i+hn},$ for some $i\in\{0,1,\ldots,d_{0}-1\}$, $j\in\{0,1,\ldots,d_{\ell}-1\}$ and $h\in\{0,\ldots,d-1\}$.
Proof.
  1. 1.
    If $d=1$, then the cyclic group $H$ generated by $(\xi^{n},\xi^{n-\ell})$ is isomorphic to the group $\langle\xi^{n}\rangle\times\langle\xi^{n-\ell}\rangle$ and has order $\frac{(q-1)^{2}}{d_{0}d_{\ell}}$. By Theorem 3.3, the number of $(n,\ell)$-equivalence classes is $d_{0}d_{\ell}$. Therefore, we can partition $\mathbb{F}_{q}^{*}\times\mathbb{F}_{q}^{*}$ as
    $\mathbb{F}_{q}^{*}\times\mathbb{F}_{q}^{*}=\bigcup_{i=0}^{d_{0}-1}\bigcup_{j=0}^{d_{\ell}-1}(\xi^{i},\xi^{j})H.$
    Then any pair $(a_{0},a_{\ell})$ is $(n,\ell)-$equivalent to one of the pairs $(\xi^{i},\xi^{j}),$ for $i=0,1,\ldots,d_{0}-1$, and $j=0,1,\ldots,d_{\ell}-1$.
  2. 2.
    If $d\neq 1,$ the number of $(n,\ell)$-equivalence classes is $dd_{0}d_{\ell},$ and so we partition $\mathbb{F}_{q}^{*}\times\mathbb{F}_{q}^{*}$ as
    $\mathbb{F}_{q}^{*}\times\mathbb{F}_{q}^{*}=\bigcup_{h=0}^{d-1}\bigcup_{i=0}^{d_{0}-1}\bigcup_{j=0}^{d_{\ell}-1}(\xi^{i+hn},\xi^{j})H,$
    which implies the second statement, similarly to the first case.
In the following, we deduce additional properties of the $(n,\ell)$-equivalence between families of $\ell$-trinomial codes, linking this $(n,\ell)$-equivalence to that of constacyclic codes from [4, 5].
Corollary 3.8.
Let $(a_{0},a_{\ell})$ and $(b_{0},b_{\ell})$ be elements of $\mathbb{F}_{q}^{*}\times\mathbb{F}_{q}^{*}$ such that $(a_{0},a_{\ell})\sim_{(n,\ell)}(b_{0},b_{\ell})$. Then, for each $i\in\{0,\ell\}$,
  1. 1.
    $a_{i}^{-1}b_{i}\in\langle\xi^{n-i}\rangle$, where $\xi$ is a primitive element of $\mathbb{F}_{q}^{*}$,
  2. 2.
    $(a_{i}^{-1}b_{i})^{d_{i}}=1$, where $d_{i}=\frac{q-1}{\gcd(n-i,q-1)}$,
  3. 3.
    $a_{i}\sim_{n-i}b_{i}$, i.e., the class of $a_{i}$-constacyclic codes of length $n-i$ is equivalent to the class of $b_{i}$-constacyclic codes of length $n-i$ over $\mathbb{F}_{q}.$
This last result implies that we can use known results about the equivalence of families of constacyclic codes for the equivalence of trinomial codes.
4 Equivalence of $p^{\ell}$-trinomial Codes of Length $n=p^{\ell+r}$.
In this section we study $p^{\ell}$-trinomial codes of length $n=p^{\ell+r}$, where $p$ is the characteristic of the underlying field $\mathbb{F}_{q}$ and $r$ an integer. First, we recall the following lemma, which combines Artin-Schreier’s theorem [Roman1995, Theorem 12.2.1] and [9, Corollary 3.79].
Lemma 4.1.
Let $a\in\mathbb{F}_{q}$ and let $p$ be the characteristic of $\mathbb{F}_{q}$. Then the trinomial $x^{p}-x-a$ is irreducible in $\mathbb{F}_{q}[x]$ if and only if $\operatorname{Tr}_{\mathbb{F}_{q}/\mathbb{F}_{p}}(a)\neq 0$; otherwise it splits (into linear factors) in $\mathbb{F}_{q}$. If it splits, then the roots are of the form $\beta+i$ for $i=0,1,\dots,p-1$ and any root $\beta$.
From the above lemma we easily deduce the following proposition.
Proposition 4.2.
Let $\mathbb{F}_{q}$ be a finite field with $q=p^{s}$ elements. Then,
  1. 1.
    The trinomial $x^{p}-x-1$ is irreducible over $\mathbb{F}_{q}$ if and only if $\gcd(s,p)=1;$ otherwise, it splits in $\mathbb{F}_{q}$.
  2. 2.
    For any $b\in\mathbb{F}_{q}^{*},$ the trinomial $b^{p}x^{p}-bx-1$ is irreducible over $\mathbb{F}_{q}$ if and only if $\gcd(s,p)=1;$ otherwise, it splits in $\mathbb{F}_{q}$.
  3. 3.
    In particular, if $p=2,$ the polynomial $x^{2}-x-1$ is irreducible over $\mathbb{F}_{2^{2k+1}}$ for any positive integer $k;$ otherwise, it splits in $\mathbb{F}_{2^{2k+1}}.$
  4. 4.
    If $p\neq 2$ then $x^{p}-x-1$ is irreducible over $\mathbb{F}_{2^{2k}}$ for any positive integer $k;$ otherwise, it splits in $\mathbb{F}_{2^{2k}}$.
Proof.
  1. 1.
    As $\operatorname{Tr}_{\mathbb{F}_{q}/\mathbb{F}_{p}}(1)=1+1+\ldots+1=s$ we have that $\operatorname{Tr}_{\mathbb{F}_{q}/\mathbb{F}_{p}}(1)=0$ in $\mathbb{F}_{p}$ if and only if $s$ is a multiple of $p$. As $p$ is a prime number, we get that $\operatorname{Tr}_{\mathbb{F}_{q}/\mathbb{F}_{p}}(1)=0$ is equivalent to $\gcd(s,p)\neq 1$. The statement now follows from Lemma 4.1.
  2. 2.
    Follows from the fact that if $f(x)$ is irreducible, then $f(bx)$ is also irreducible, for each $b\in\mathbb{F}_{q}^{*}$, see [9, p. 121].
  3. 3.
    and 4. are direct applications of 1.
Corollary 4.3.
For each integer $\ell,$ the polynomial $x^{p^{\ell+1}}-x^{p^{\ell}}-1$ can be factorized into irreducible factors over $\mathbb{F}_{q}$ (where $q=p^{s}$) as follows:
$x^{p^{\ell+1}}-x^{p^{\ell}}-1=\begin{cases}\left(x^{p}-x-1\right)^{p^{\ell}},&\text{if $\gcd(s,p)=1$ }\\ \displaystyle{\prod_{i=0}^{p-1}}(x-(\beta+i))^{p^{\ell}},&\text{else }\\ \end{cases}$
where $\beta\in\mathbb{F}_{q}$ is a root of $x^{p}-x-1.$
Proof.
As the characteristic of $\mathbb{F}_{q}$ is $p,$ we have by Proposition 4.2 that
$x^{p^{\ell+1}}-x^{p^{\ell}}-1=\left(x^{p}-x-1\right)^{p^{\ell}}=\begin{cases}\left(x^{p}-x-1\right)^{p^{\ell}},&\text{if $\gcd(s,p)=1$ }\\ \displaystyle{\prod_{i=0}^{p-1}}(x-(\beta+i))^{p^{\ell}},&\text{else }\\ \end{cases}$
where $\beta$ is a root of $x^{p}-x-1.$
We can now derive results about the possible generator polynomials of codes associated to some specific trinomials. We start with codes of length $n=p^{\ell+1}$:
Theorem 4.4.
Let $\ell$ be an integer and $a,b$ be two elements of $\mathbb{F}_{q}^{*}$ such that $a^{p}=b^{p-1}$. Then each $p^{\ell}$-trinomial code associated with the polynomial $x^{p^{\ell+1}}-ax^{p^{\ell}}-b$ has a polynomial generator of the form
$g(x)=\begin{cases}\left((\alpha x)^{p}-\alpha x-1\right)^{j},&\text{if $\gcd(s,p)=1$ }\\ \displaystyle{\prod_{i=0}^{p-1}}(\alpha x-(\beta+i))^{j},&\text{else }\\ \end{cases}\quad\text{with }1\leq j\leq p^{\ell},$
where $\alpha=(ba^{-1})^{p^{-\ell}}$, and $\beta$ a root of $x^{p}-x-1.$
Proof.
As $a^{p}=b^{p-1}$, we have from Lemma 2.7 that
$\gcd(x^{p^{\ell+1}}-b,x^{p^{\ell+1}-p^{\ell}}-a)=x^{\gcd(p^{\ell+1}-p^{\ell},p^{\ell})}-c=x^{p^{\ell}}-c=(x-\alpha)^{p^{\ell}},$
for some $c$ and $\alpha$ in $\mathbb{F}_{q}^{*}$ such that $c=\alpha^{p^{\ell}}$. This polynomial has a root $\alpha$ with multiplicity $p^{\ell}$. Since $\alpha$ is a common root of $x^{p^{\ell+1}}-b$ and $x^{p^{\ell+1}-p^{\ell}}-a$, it follows that $\alpha^{p^{\ell+1}}=b$ and $\alpha^{p^{\ell+1}-p^{\ell}}=a$. Thus,
$ba^{-1}=\frac{\alpha^{p^{\ell+1}}}{\alpha^{p^{\ell+1}-p^{\ell}}}=\alpha^{p^{\ell}}.$
Now we can factor $x^{p^{\ell+1}}-ax^{p^{\ell}}-b$ as follows:
$\begin{array}[]{rl}x^{p^{\ell+1}}-ax^{p^{\ell}}-b&=b\left(b^{-1}x^{p^{\ell+1}}-ab^{-1}x^{p^{\ell}}-1\right)\\ &=b\left(\alpha^{-p^{\ell+1}}x^{p^{\ell+1}}-\alpha^{-p^{\ell}}x^{p^{\ell}}-1\right)\\ &=b\left((\alpha^{-1}x)^{p^{\ell+1}}-(\alpha^{-1}x)^{p^{\ell}}-1\right)\\ &=b\left((\alpha^{-1}x)^{p}-\alpha^{-1}x-1\right)^{p^{\ell}}.\end{array}$
According to Lemma 4.1, $x^{p}-x-1$ is irreducible because $\gcd(s,p)=1;$ otherwise, it splits in $\mathbb{F}_{q}$. Therefore, $x^{p^{\ell+1}}-ax^{p^{\ell}}-b$ can be factored over $\mathbb{F}_{q}$ as follows:
$x^{p^{\ell+1}}-ax^{p^{\ell}}-b=\begin{cases}b\left((\alpha^{-1}x)^{p}-(\alpha^{-1}x)-1\right)^{p^{\ell}},&\text{if $\gcd(s,p)=1$ }\\ \displaystyle{\prod_{i=0}^{p-1}}(\alpha x-(\beta+i))^{p^{\ell}},&\text{else }\end{cases}.$
Now any $p^{\ell}$-trinomial code associated with the polynomial $x^{p^{\ell+1}}-ax^{p^{\ell}}-b$ has a generator polynomial of the desired form.
We then turn to codes of length $n=p^{\ell+s}$, where $s$ is the extension degree of $\mathbb{F}_{q}$. For this we first need the following lemma.
Lemma 4.5.
[9, Theorem. 3.80.] For $x^{q}-x-a$ with $a$ an element of the subfield $\mathbb{F}_{p^{r}}$ of $\mathbb{F}_{q},\ q=p^{s}$, we have the decomposition
$x^{q}-x-a=\prod_{j=1}^{q/p^{r}}\left(x^{p^{r}}-x-\beta_{j}\right)=\prod_{j=1}^{p^{s-r}}\left(x^{p^{r}}-x-\beta_{j}\right)$
in $\mathbb{F}_{q}[x]$, where the $\beta_{j}$ are the distinct elements of $\mathbb{F}_{q}$ with $\operatorname{Tr}_{\mathbb{F}_{q}/\mathbb{F}_{p^{r}}}\left(\beta_{j}\right)=a$.
Theorem 4.6.
Let $\mathbb{F}_{q}$ be a finite field with $q=p^{s}$ elements, and $\ell$ a positive integer. Then the polynomial $x^{p^{\ell+s}}-x^{p^{\ell}}-1$ has the following irreducible decomposition:
$x^{p^{\ell+s}}-x^{p^{\ell}}-1=\prod_{j=1}^{p^{s-1}}\left(x^{p}-x-\beta_{j}\right)^{p^{\ell}},$
where $\operatorname{Tr}_{\mathbb{F}_{q}/\mathbb{F}_{p}}(\beta_{j})=1$.
Proof.
Since the characteristic of $\mathbb{F}_{q}$ is $p$, we have
$x^{p^{\ell+s}}-x^{p^{\ell}}-1=\left(x^{p^{s}}-x-1\right)^{p^{\ell}}=\left(x^{q}-x-1\right)^{p^{\ell}}.$
As $1\in\mathbb{F}_{p}$, by Lemma 4.5 we obtain
$x^{p^{\ell+s}}-x^{p^{\ell}}-1=\prod_{j=1}^{p^{s-1}}\left(x^{p}-x-\beta_{j}\right)^{p^{\ell}},$
where $\operatorname{Tr}_{\mathbb{F}_{q}/\mathbb{F}_{p}}(\beta_{j})=1$. According to Lemma 4.1, since $\operatorname{Tr}_{\mathbb{F}_{q}/\mathbb{F}_{p}}(\beta_{j})=1\neq 0$, the polynomial $x^{p}-x-\beta_{j}$ is irreducible over $\mathbb{F}_{q}$ for each $j=1,\ldots,p^{s-1}$.
Corollary 4.7.
Let $\ell$ be an integer and $a,b$ be two elements of $\mathbb{F}_{q}^{*}$, where $\ q=p^{s}$. If $(a,b)\sim_{(p^{\ell+s},p^{\ell})}(1,1)$ then $a=1$ and each $p^{\ell}$-trinomial code $C$ associated with the polynomial $x^{p^{\ell+s}}-x^{p^{\ell}}-b,$ with $b\in\mathbb{F}_{q}^{*},$ has a polynomial generator of the form
$g(x)=\prod_{j=1}^{p^{s-1}}\left((\alpha^{-1}x)^{p}-\alpha^{-1}x-\beta_{j}\right)^{i},\ \text{with}\ 0\leq i\leq p^{\ell},$
for some $\beta_{j}\in\mathbb{F}_{q}^{*}$ with $\operatorname{Tr}_{\mathbb{F}_{q}/\mathbb{F}_{p}}(\beta_{j})=1$ and $\alpha\in\mathbb{F}_{q}^{*}$ such that $\alpha=b^{-p^{-\ell}}.$  [^2] Such an $\alpha$ always exists since the equation $x^{p^{\ell}}=b^{-1}$ has a unique solution in $\mathbb{F}_{q}$.
Proof.
Suppose that $(a,b)\sim_{(p^{\ell+s},p^{\ell})}(1,1)$ then by Corollary 3.5 there exists an $\alpha\in\mathbb{F}_{q}^{*}$ such that
$(b\alpha^{p^{s+\ell}},a\alpha^{p^{s+k}-p^{\ell}})=(b\alpha^{qp^{\ell}},a\alpha^{(p^{s}-1)p^{\ell}})=(b\alpha^{p^{\ell}},a\alpha^{(q-1)^{p^{\ell}}})=(b\alpha^{p^{\ell}},a)=(1,1).$
Thus $a=1$ and $\alpha=b^{-p^{-\ell}}$. The rest of the proof is a direct application of Theorem 4.6.
5 Examples of Families of Trinomial Codes
In this section, we give some examples of restricting the search space for good trinomial codes by applying the theory developed before.
Example 5.1 ($3^{\ell}$-trinomial codes of length $n=3^{\ell+1}$ over $\mathbb{F}_{3}$ ).
We consider $3^{\ell}$-trinomial codes of length $n=3^{\ell+1}$ over $\mathbb{F}_{3}.$ By Theorem 3.3, the number of $(3^{\ell+1},3^{\ell})$-equivalence classes is equal to
$N=\frac{4}{\operatorname{lcm}(\frac{2}{\gcd(3^{\ell+1},2)},\frac{2}{\gcd(3^{\ell+1}-3^{\ell},2)})}=\frac{4}{\operatorname{lcm}(2,\frac{2}{\gcd(2.3^{\ell},2)})}=2.$
As $\gcd(\frac{2}{\gcd(3^{\ell+1},2)},\frac{2}{\gcd(3^{\ell+1}-3^{\ell},2)})=(2,1)=1$ – according to Theorem 3.7 – each $3^{\ell}$-trinomial codes of length $n=3^{\ell+1}$ is equivalent to a $3^{\ell}$-trinomial code associated with $x^{3^{\ell+1}}-x^{3^{\ell}}-1$ or $x^{3^{\ell+1}}-2x^{3^{\ell}}-1.$
According to Corollary 3.5, $(a_{0},a_{\ell})\sim_{(3^{\ell+1},3^{\ell})}(1,1)$ if there is $\alpha\in\mathbb{F}_{3}^{*}$ such that
$(a_{0},a_{\ell})\star(\alpha^{3^{\ell+1}},\alpha^{3^{\ell+1}-3^{\ell}})=(a_{0}\alpha^{3^{\ell+1}},a_{\ell}\alpha^{2\cdot 3^{\ell}})=(a_{0}\alpha,a_{\ell})\stackrel{{\scriptstyle!}}{{=}}(1,1).$
Thus $a_{\ell}=1$ and $a_{0}=\alpha^{-1}$ for $\alpha=1,2$. Therefore, the equivalence class of $(1,1)$ consists of the pairs $(1,1)$ and $(2,1)$. Similarly the class of $(1,2)$ consists of the pairs $(1,2)$ and $(2,2).$
Since $\gcd(3,1)=1,$ by Corollary 4.3 the factorization of $x^{3^{\ell+1}}-x^{3^{\ell}}-1$ is given by
$x^{3^{\ell+1}}-x^{3^{\ell}}-1=(x^{3}-x-1)^{3^{\ell}}.$
According to Theorem 4.4, each code in the class of $(1,1)$ has a generator polynomial of the form $g(x)=(x^{3}-x-1)^{i}$, for some $0\leq i\leq 3^{\ell}$. This class contains some optimal codes, for example, by taking $\ell=2,$ we found that the $9$-trinomial code associated with $x^{27}+2x^{9}+2$ and generated by $g(x)=x^{24}+x^{22}+x^{21}+x^{20}+2x^{19}+2x^{18}+x^{16}+2x^{15}+x^{14}+x^{11}+x^{9}+x^{8}+x^{7}+x^{6}+2x^{5}+2x^{4}+x^{2}+2x+1$ is an optimal $[27,3,18]_{3}$-code. [^3] These codes attain the Griesmer bound.
For $\ell=3,$ we found that the $27$-trinomial code associated with $x^{27}+2x^{9}+2$ and generated by $g(x)=x^{78}+x^{76}+x^{75}+x^{74}+2x^{73}+2x^{72}+x^{70}+2x^{69}+x^{68}+x^{65}+x^{63}+x^{62}+x^{61}+2x^{60}+2x^{59}+x^{57}+2x^{56}+x^{55}+x^{52}+x^{50}+x^{49}+x^{48}+2x^{47}+2x^{46}+x^{44}+2x^{43}+x^{42}+x^{39}+x^{37}+x^{36}+x^{35}+2x^{34}+2x^{33}+x^{31}+2x^{30}+x^{29}+x^{26}+x^{23}+x^{21}+x^{20}+x^{19}+2x^{18}+2x^{17}+x^{15}+2x^{14}+x^{13}+x^{10}+x^{8}+x^{7}+x^{6}+2x^{5}+2x^{4}+x^{2}+2x+1$ is an optimal $[81,3,55]_{3}$-code.33footnotemark: 3
Example 5.2 ($3^{\ell}$-trinomial codes of length $n=3^{\ell+2}$ over $\mathbb{F}_{9}$).
We consider $3^{\ell}$-trinomial codes of length $n=3^{\ell+2}$ over $\mathbb{F}_{9}=\mathbb{F}_{3}(\xi)$ with $\xi$ a primitive element of $\mathbb{F}_{9}$. The number of $(3^{\ell+2},3^{\ell})$-equivalence classes is
$N=\frac{64}{\operatorname{lcm}(\frac{8}{\gcd(3^{\ell+2},8)},\frac{8}{\gcd(3^{\ell+2}-3^{\ell},8)})}=\frac{64}{\operatorname{lcm}(8,\frac{8}{\gcd(8.3^{\ell},8)})}=\frac{64}{\operatorname{lcm}(8,1)}=8,$
and each class contains $8$ pairs. Since $\gcd(\frac{8}{\gcd(3^{\ell+2},8)},\frac{8}{\gcd(8.3^{\ell},8)})=\gcd(8,1)=1,$ then by Theorem 3.7, the $8$ possible pairs are
$(1,1),(1,\xi),(1,\xi^{2}),(1,\xi^{3}),(1,\xi^{4}),(1,\xi^{5}),(1,\xi^{6}),(1,\xi^{7}).$
It follows that for $a,b\in\mathbb{F}_{9}^{*}$ the $3^{\ell}$-trinomial code family associated with $x^{3^{\ell+2}}-ax^{3^{\ell}}-b,$ is equivalent to a $3^{\ell}$-trinomial code family associated with one of the polynomials $x^{3^{\ell+2}}-x^{3^{\ell}}-\xi^{j},\ j=0,1,\ldots,7.$
We now determine the class of $(1,1)$. According to Corollary 3.5, $(a_{0},a_{\ell})\sim_{(3^{\ell+2},3^{\ell})}(1,1)$ if there exists $\alpha\in\mathbb{F}_{9}^{*}$ such that
$(a_{0},a_{\ell})\star(\alpha^{3^{\ell+2}},\alpha^{3^{\ell+2}-3^{\ell}})=(a_{0}\alpha^{3^{\ell+2}},a_{\ell}\alpha^{8\cdot 3^{\ell}})=(a_{0}\alpha^{3^{\ell}},a_{\ell})=(1,1).$
Thus, $a_{\ell}=1$ and $a_{0}=\alpha^{-3^{\ell}}$. Therefore, the class of $(1,1)$ is given by
$\{(\alpha^{-3^{\ell}},1):\alpha\in\mathbb{F}_{9}^{*}\}=\{(1,1),(\xi,1),(\xi^{2},1),(\xi^{3},1),(\xi^{4},1),(\xi^{5},1),(\xi^{6},1),(\xi^{7},1)\}.$
Note that the second equality follows from the fact that the order of $\mathbb{F}_{9}^{*}$ is $8$ and thus coprime to $3^{\ell}$ for any $\ell$, which implies that all elements of $\mathbb{F}_{9}^{*}$ appear in the first coordinate.
According to Theorem 4.6, the factorization of $x^{3^{\ell+2}}-x^{3^{\ell}}-1$ is given by:
$x^{3^{\ell+2}}-x^{3^{\ell}}-1=(x^{9}-x-1)^{3^{\ell}}=\prod_{j=1}^{3}\left(x^{3}-x-\beta_{j}\right)^{3^{\ell}},$
for some $\beta\in\mathbb{F}_{q}^{*}$ with $\operatorname{Tr}_{\mathbb{F}_{9}/\mathbb{F}_{3}}(\beta_{j})=1$. It follows that
$x^{3^{\ell+2}}-x^{3^{\ell}}-1=\left(x^{3}-x-\xi\right)^{3^{\ell}}\left(x^{3}-x-2\right)^{3^{\ell}}\left(x^{3}-x-\xi^{3}\right)^{3^{\ell}}$
Hence, the $3^{\ell}$-trinomial codes of length $n=3^{\ell+2}$ over $\mathbb{F}_{9}$ which are equivalent to $3^{\ell}$-trinomial codes associated with $x^{3^{\ell+2}}-x^{3^{\ell}}-1$ will have a generator polynomial of the form
$g(x)=\left(x^{3}-x-\xi\right)^{i}\left(x^{3}-x-2\right)^{j}\left(x^{3}-x-\xi^{3}\right)^{h},\quad 0\leq i,j,h\leq 3^{\ell}.$
For $\ell=1,$ we constructed the $3-$trinomial code associated to $x^{27}-x^{3}-1$, and generated by $g(x)=x^{3}-x+\xi^{7}$ which is an optimal $[27,24,3]_{9}$-code. [^4] These codes attain an upper bound on the minimum distance according to [8].
For $\ell=2,$ we constructed the $9-$trinomial code associated to $x^{81}-x^{9}-1$, and generated by $g(x)=x^{3}-x+\xi^{5}$ which is an optimal $[81,78,3]_{9}$-code.44footnotemark: 4
Example 5.3 (Trinomial codes of length $27$ over $\mathbb{F}_{4}$).
We consider the case of $\ell$-trinomial codes of length $27$ over $\mathbb{F}_{4}=\mathbb{F}_{2}(\xi),$ with $\xi$ a primitive element of $\mathbb{F}_{4}$. The number of $(27,\ell)$-equivalence classes is given by
$\frac{3^{2}}{\operatorname{lcm}\left(\frac{3}{\gcd(27,3)},\frac{3}{\gcd(27-\ell,3)}\right)}=\frac{9}{\operatorname{lcm}\left(1,\frac{3}{\gcd(27-\ell,3)}\right)}.$
We have two cases:
  1. 1.
    If $k\equiv 0\pmod{3}$: The number of $(27,\ell)$-equivalence classes is $9$. In this case, we consider all pairs $(a,b)$ from $\mathbb{F}_{4}^{*}\times\mathbb{F}_{4}^{*}$.
  2. 2.
    If $\ell\not\equiv 0\pmod{3}$: The number of $(27,\ell)$-equivalence classes is $3$. Since
    $\gcd\left(\frac{3}{\gcd(27,3)},\frac{3}{\gcd(27-\ell,3)}\right)=\gcd(1,3)=1,$
    then – by Theorem 3.7 – the representatives of these three $(n,\ell)$-equivalence classes are the pairs $(1,1)$, $(1,\xi)$, and $(1,\xi^{2})$. So we need to consider the three polynomials $x^{27}-x^{\ell}-1,\ \ x^{27}-x^{\ell}-\xi,$ and $x^{27}-x^{\ell}-\xi^{2}.$
    For $\ell=8,$ we found that the $8$-trinomial code associated to $x^{27}-x^{8}-\xi^{2}$ and generated by $x^{6}+\xi^{2}x^{5}+\xi^{2}x^{3}+\xi x^{2}+x+1$ is a $[27,21,4]_{4}$-code, which equals the best known parameters according to [8].
    For $\ell=5,$ the $5$-trinomial code associated with $x^{27}-x^{5}-1$ and generated by $x^{10}+x^{9}+x^{8}+x^{7}+x^{6}+\xi^{2}x^{5}+x^{4}+\xi^{2}x^{2}+\xi x+1$ is a $[27,17,6]_{4}$-code, which equals the best known parameters according to [8].
6 Equivalence of Polycyclic Codes
In this section we generalize the results on equivalence to general polycyclic codes. We start with the generalized definition of equivalence, for which we will use the notation $\vec{a}(x):=a_{0}+a_{1}x+\ldots+a_{n-1}x^{n-1}$.
Definition 6.1.
Let $\vec{a}=(a_{0},a_{1},\ldots,a_{n-1})$ and $\vec{b}=(b_{0},b_{1},\ldots,b_{n-1})$ be elements in $\mathbb{F}_{q}^{n}.$ We say that $\vec{a}$ and $\vec{b}$ are $n$-equivalent, and we denote this by
$\vec{a}\sim_{n}\vec{b},$
if there exists an $\alpha\in\mathbb{F}_{q}^{*}$ such that the following map
$\begin{array}[]{cccc}\varphi_{\alpha}:&\mathbb{F}_{q}[x]/\langle x^{n}-\vec{b}(x)\rangle&\longrightarrow&\mathbb{F}_{q}[x]/\langle x^{n}-\vec{a}(x)\rangle,\\ &f(x)&\longmapsto&f(\alpha x),\end{array}$ (6)
is an $\mathbb{F}_{q}$-algebra isomorphism.
Note that, as before, the map $\varphi_{\alpha}$ is a Hamming isometry. Moreover, we can easily verify that "$\sim_{n}"$ is an equivalence relation.
Remark 6.2.
  1. 1.
    If $\vec{a}=(\lambda,0,\ldots,0)$ and $\vec{b}=(\mu,0,\ldots,0)$, then we recover the case of $n$-equivalence for constacyclic codes studied in [4].
  2. 2.
    If $\vec{a}=(a_{0},0,\ldots,a_{\ell},0,\ldots,0)$ and $\vec{b}=(b_{0},0,\ldots,b_{\ell},0,\ldots,0)$, then we recover the case of $(n,\ell)$-equivalence for $\ell$-trinomial codes studied in Section 3.
In the following we show that this notion of equivalence automatically implies that the vectors $\vec{a}$ and $\vec{b}$ have zero entries in exactly the same position. This implies that $\ell$-trinomial code families can only be equivalent to other $\ell$-trinomial code families.
Lemma 6.3.
Let $\vec{a}=(a_{0},a_{1},\ldots,a_{n-1})$ and $\vec{b}=(b_{0},b_{1},\ldots,b_{n-1})$ be elements in $\mathbb{F}_{q}^{n}$ such that $\vec{a}\sim_{n}\vec{b}$. Then $a_{i}\neq 0$ if and only if $b_{i}\neq 0$, for any $0\leq i\leq n-1,$ and so $\vec{a}$ and $\vec{b}$ have the same Hamming weight.
Proof.
Suppose that $\vec{a}\sim_{n}\vec{b},$ then there is $\alpha\in\mathbb{F}_{q}^{*}$ such that $\varphi_{\alpha}$ is an $\mathbb{F}_{q}$-algebra isometry between $\mathbb{F}_{q}[x]/\langle x^{n}-\vec{b}(x)\rangle$ and $\mathbb{F}_{q}[x]/\langle x^{n}-\vec{a}(x)\rangle$. Then, as in the proof of [Theorem 3.3, (1) $\Rightarrow$ (2)], we obtain that
$b_{i}=\alpha^{n-i}a_{i},\quad\forall i=0,\ldots,n-1.$
Hence the result holds.
We now generalize Theorem 3.3 to the general polycyclic case:
Theorem 6.4.
Let $\vec{a}=(a_{0},a_{1},\ldots,a_{n-1}),\vec{b}=(b_{0},b_{1},\ldots,b_{n-1})\in\mathbb{F}_{q}^{n}$ have non-zero entries in the same $m$ positions, i.e., $a_{i_{j}}$ and $b_{i_{j}}$ are non-zero for $10\leq i_{0}<\dots<i_{m-1}\leq n-$ 1. Moreover, let $\xi$ be a primitive element of $\mathbb{F}_{q}$. Then the following statements are equivalent:
  1. 1.
    $\vec{a}\sim_{n}\vec{b}$.
  2. 2.
    The polynomials $a_{i_{j}}x^{n-i_{j}}-b_{i_{j}}\in\mathbb{F}_{q}[x]$, for $j\in\{0,1,\ldots,m-1\}$, have a common root in $\mathbb{F}_{q}^{*}$.
  3. 3.
    The polynomial $\gcd(a_{i_{0}}x^{n-i_{0}}-b_{0},a_{i_{1}}x^{n-i_{1}}-b_{i_{1}},\ldots,a_{i_{m-1}}x^{n-i_{m-1}}-b_{i_{m-1}})$ has at least one root in $\mathbb{F}_{q}^{*}$.
  4. 4.
    The polynomial $\gcd_{\{0\leq j\leq m-1\}}(x^{n-{i_{j}}}-b_{i_{j}}a_{i_{j}}^{-1})$ has at least one root in $\mathbb{F}_{q}^{*}$.
  5. 5.
    There exists $\alpha\in\mathbb{F}_{q}^{*}$ such that
    $(a_{i_{0}},a_{i_{1}},\ldots,a_{i_{m-1}})\star(\alpha^{n-{i_{0}}},\alpha^{n-{i_{1}}},\ldots,\alpha^{n-{i_{m-1}}})=(b_{i_{0}},b_{i_{1}},\ldots,b_{i_{m-1}}).$
  6. 6.
    $(a_{i_{0}},a_{i_{1}},\ldots,a_{i_{m-1}})^{-1}\star(b_{i_{0}},b_{i_{1}},\ldots,b_{i_{m-1}})\in H,$ where $H$ is the cyclic subgroup of $(\mathbb{F}_{q}^{*})^{m}$ generated by $(\xi^{n-{i_{0}}},\xi^{n-{i_{1}}},\ldots,\xi^{n-{i_{m-1}}}).$
In particular the number of $n$-equivalence classes is
$N=\dfrac{(q-1)^{m}}{\operatorname{lcm}_{(0\leq j\leq m-1)}\left(\frac{q-1}{\gcd(n-i_{j},q-1)}\right)}.$
Proof.
  1. (1) $\Rightarrow$ (2)
    Suppose that $\vec{a}\sim_{n}\vec{b}$, then there is $\alpha\in\mathbb{F}_{q}^{*}$ such that
    $\varphi_{\alpha}:\mathbb{F}_{q}[x]/\langle x^{n}-\vec{b}(x)\rangle\to\mathbb{F}_{q}[x]/\langle x^{n}-\vec{a}(x)\rangle,\quad f(x)\mapsto f(\alpha x)$
    is an $\mathbb{F}_{q}$-algebra isometry. It follows that
    $\varphi_{\alpha}(x^{k})=\alpha^{k}x^{k}\mod(x^{n}-\vec{a}(x)),\quad\forall\ k=0,1,\ldots,n-1.$
    As $\varphi_{\alpha}$ is an $\mathbb{F}_{q}$-algebra isometry and $\varphi(x^{n}-\vec{b}(x))=0\mod(x^{n}-\vec{a}(x))$, then
    $\varphi_{\alpha}(x^{n})=\varphi_{\alpha}(\vec{b}(x))=b_{0}+\alpha b_{1}x+\ldots+\alpha^{n-1}b_{n-1}x^{n-1},\mod(x^{n}-\vec{a}(x)).$ (7)
    On the other hand,
    $\varphi_{\alpha}(x^{n})=\alpha^{n}x^{n}=\alpha^{n}(a_{0}+a_{1}x+\ldots+a_{n-1}x^{n-1}),\mod(x^{n}-\vec{a}(x))$ (8)
    Comparing term by term, we deduce that for any $i\in\{0,1,\ldots,n-1\},\ a_{i}\alpha^{n-i}=b_{i}$, which means that $\alpha$ is a common root of the polynomials $a_{i}x^{n-i}-b_{i}$, for $i\in\{0,1,\ldots,n-1\}$. As $a_{i_{j}}$’s and $b_{i_{j}}$’s are the non-zeros components of $\vec{a}$ and $\vec{b},$ then $a_{i_{j}}\alpha^{n-{i_{j}}}=b_{i_{j}},\ j=0,\ldots,m-1$.
  2. (2) $\Rightarrow$ (3)
    and (3) $\Rightarrow$ (4) are immediate.
  3. (4) $\Rightarrow$ (5)
    Let $\alpha$ be a root of the polynomial $\gcd_{\{0\leq i\leq n-1,\ a_{i}\neq 0\}}(x^{n-i}-b_{i}a_{i}^{-1})$. Then $\alpha$ is a common root of the polynomials $x^{n-i_{j}}-b_{i_{j}}a_{i_{j}}^{-1}\ \text{for any}\ j\in\{0,1,\ldots,m-1\},$ and so $a_{i_{j}}\alpha^{n-i_{j}}=b_{i_{j}}$. It follows that
    $(b_{i_{0}},b_{i_{1}},\ldots,b_{i_{m-1}})=(\alpha^{n-i_{0}},\alpha^{n-i_{1}},\ldots,\alpha^{n-i_{m-1}})\star(a_{i_{0}},a_{i_{1}},\ldots,a_{i_{m-1}}).$
  4. (5) $\Rightarrow$ (6)
    Suppose that there is $\alpha\in\mathbb{F}_{q}^{*}$ such that
    $(b_{i_{0}},b_{i_{1}},\ldots,b_{i_{m-1}})=(\alpha^{n-i_{0}},\alpha^{n-i_{1}},\ldots,\alpha^{n-i_{m-1}})\star(a_{i_{0}},a_{i_{1}},\ldots,a_{i_{m-1}}).$
    For $\alpha=\xi^{h},$ we obtain that
    $(a_{i_{0}},a_{i_{1}},\ldots,a_{i_{m-1}})^{-1}\star(b_{i_{0}},b_{i_{1}},\ldots,b_{i_{m-1}})=(\xi^{n-{i_{0}}},\xi^{n-{i_{1}}},\ldots,\xi^{n-i_{m-1}})^{h}.$
    It follows that $(a_{i_{0}},a_{i_{1}},\ldots,a_{i_{m-1}})^{-1}\star(b_{i_{0}},b_{i_{1}},\ldots,b_{i_{m-1}})$ belongs to the cyclic subgroup $H$ of $(\mathbb{F}_{q}^{*})^{m}$ generated by $(\xi^{n-{i_{0}}},\xi^{n-{i_{1}}},\ldots,\xi^{n-{i_{m-1}}}).$
  5. (6) $\Rightarrow$ (1)
    Suppose that $(a^{n-{i_{0}}},a^{n-{i_{1}}},\ldots,a^{n-i_{m-1}})^{-1}\star(b^{n-{i_{0}}},b^{n-{i_{1}}},\ldots,b^{n-i_{m-1}})$ is an element of the cyclic subgroup $H$ of $(\mathbb{F}_{q}^{*})^{m}$ generated by $(\xi^{n-{i_{0}}},\xi^{n-{i_{1}}},\ldots,\xi^{n-{i_{m-1}}})$. Then there exists an integer $h$ such that
    $(a_{i_{0}},a_{i_{1}},\ldots,a_{i_{m-1}})^{-1}\star(b_{i_{0}},b_{i_{1}},\ldots,b_{i_{m-1}})=(\xi^{n-{i_{0}}},\xi^{n-{i_{1}}},\ldots,\xi^{n-{i_{m-1}}})^{h}$
    For $\beta=\xi^{h},$ we obtain that $a_{j}\beta^{n-j}=b_{j},$ for any $j\in\{i_{0},i_{1},\ldots,i_{m-1}\}.$ As in the proof of Theorem 3.3, we verify that $\varphi_{\beta},$ as follows:
    $\begin{array}[]{cccc}\varphi_{\beta}:&\mathbb{F}_{q}[x]/\langle x^{n}-\vec{b}(x)\rangle,&\longrightarrow&\mathbb{F}_{q}[x]/\langle x^{n}-\vec{a}(x)\rangle,\\ &f(x)&\longmapsto&f(\beta x),\end{array}$ (9)
    is an $\mathbb{F}_{q}$-algebra isometry with respect to the Hamming distance.
By the equivalence between (1) and (6), we deduce that the number of $n$-equivalence classes on on $(\mathbb{F}_{q}^{*})^{m}$ corresponds to the order of the group $(\mathbb{F}_{q}^{*})^{m}/H,$ which equals
$N=\dfrac{(q-1)^{m}}{\operatorname{lcm}_{(0\leq j\leq m-1)}\left(\frac{q-1}{\gcd(n-i_{j},q-1)}\right)}.$
Similarly to the case of $\ell$-trinomial codes, Theorem 6.4 implies the following results regarding the equivalence of polycyclic codes. The proofs are analogous to the trinomial case.
Corollary 6.5.
As before let $\vec{a}=(a_{0},a_{1},\ldots,a_{n-1})$ and $\vec{b}=(b_{0},b_{1},\ldots,b_{n-1})$ be elements of $\mathbb{F}_{q}^{n}$ of the same weight $m$ and denote by $a_{i_{j}}$ and $b_{i_{j}}$ the non-zeros components of $\vec{a}$ and $\vec{b}$.
  1. 1.
    The class of polycyclic codes associated with the polynomial $x^{n}-\displaystyle{\sum_{j=0}^{m-1}}a_{i_{j}}x^{i_{j}}$ is equivalent to the class of polycyclic codes associated with the polynomial $x^{n}-\displaystyle{\sum_{j=0}^{m-1}}x^{i_{j}}$ if and only if there exists $\alpha\in\mathbb{F}_{q}^{*}$ such that
    $(a_{i_{0}},a_{i_{1}},\ldots,a_{i_{m-1}})\star(\alpha_{i_{0}}^{n-i_{0}},\alpha_{i_{1}}^{n-i_{1}},\ldots,\alpha_{i_{m-1}}^{n-i_{m-1}})=(1,1,\ldots,1).$
  2. 2.
    If $\gcd_{0\leq j\leq m-1}\left(\frac{q-1}{\gcd(n-i_{j},q-1)}\right)=1$ then the class of polycyclic codes associated with the polynomial $x^{n}-\displaystyle{\sum_{j=0}^{m-1}}a_{i_{j}}x^{i_{j}}$ is equivalent to the class of polycyclic codes associated with the polynomial $x^{n}-\displaystyle{\sum_{j=0}^{m-1}}\xi^{k_{j}}x^{i_{j}},$ for $k_{j}=0,1,\ldots,\gcd(n-i_{j},q-1),$ where $\xi$ is a primitive element of $\mathbb{F}_{q}$.
  3. 3.
    The number of $\mathbb{F}_{q}$-algebra isometries $\varphi_{\alpha}$ between $\mathbb{F}_{q}[x]/\langle x^{n}-\vec{b}(x)\rangle$ and ${\mathbb{F}_{q}[x]/\langle x^{n}-\vec{a}(x)\rangle}$ is
    $\gcd(n-i_{0},n-i_{1},\ldots,n-i_{m-1},q-1).$
Example 6.6 (Polycyclic codes of length $n=12$ over $\mathbb{F}_{3}$ ).
We consider polycyclic codes of length $n=12$ over $\mathbb{F}_{3}$ associated with a polynomial of the form $f(x)=x^{12}-cx^{7}-bx-a\in\mathbb{F}_{3}[x]$. Denote $\vec{a}(x):=cx^{7}+bx+a$, then the Hamming weight of $\vec{a}(x)$ is $3$ and according to Theorem 6.4, the number of $12$-equivalence classes is
$N=\frac{2^{3}}{\operatorname{lcm}(\frac{2}{\gcd(12,2)},\frac{2}{\gcd(12-1,2)},\frac{2}{\gcd(12-7,2)})}=\frac{2^{3}}{\operatorname{lcm}(1,2,2)}=4.$
Since
$\gcd(\frac{2}{\gcd(12,2)},\frac{2}{\gcd(12-1,2)},\frac{2}{\gcd(12-7,2)})=\gcd(1,2,2)=1,$
then by Corollary 6.5, each polycyclic codes associated with a polynomial $f$ of the form $f=x^{12}-cx^{7}-bx-a,$ is equivalent to a polycyclic code associated with one of the following polynomials:
$f_{1}=x^{12}-x^{7}-x-1,\quad f_{2}=x^{12}-x^{7}-x-\xi,\quad f_{3}=x^{12}-\xi x^{7}-x-1,\quad f_{4}=x^{12}-\xi x^{7}-x-\xi,$
where $\xi=2$ is (the only) primitive element of $\mathbb{F}_{3}$. We then searched for good codes in the corresponding spaces. According to Codes Tables [8] some of these codes are optimal (for given $n,k$ and $q$); we present these code parameters in Table 1.
Class polynomial $f(x)$ Generator polynomial of the polycyclic code Parameters
$f_{1}(x)=x^{12}-x^{7}-x-1$ $x^{10}+2x^{8}+x^{6}+2x^{5}+2x^{4}+x^{3}+x^{2}+2x+2$ $[12,2,9]_{3}$
$x^{8}+x^{6}+2x^{3}+2x^{2}+2x+2$ $[12,4,6]_{3}$
$f_{3}(x)=x^{12}-2x^{7}-x-1$ $x^{2}+2x+2$ $[12,10,2]_{3}$
$x^{7}+x^{6}+2x^{5}+x^{4}+2x^{2}+2$ $[12,5,6]_{3}$
$x^{5}+x^{3}+x^{2}+2x+1$ $[12,7,4]_{3}$
$x^{4}+x^{3}+2x^{2}+2$ $[12,8,3]_{3}$
$f_{4}(x)=x^{12}-2x^{7}-x-2$ $x^{3}+x^{2}+2$ $[12,9,3]_{3}$
Table 1: Optimal polycyclic codes of length $n=12$ over $\mathbb{F}_{3}$.
Example 6.7 (Polycyclic codes of length $n=15$ over $\mathbb{F}_{4}$ ).
We consider skew polycyclic codes of length $n=15$ over $\mathbb{F}_{4}=\mathbb{F}_{2}(\xi),\ \xi$ a primitive element of $\mathbb{F}_{4},$ associated with a polynomial of the form $f(x)=x^{15}-cx^{h}-bx^{l}-a\in\mathbb{F}_{4}[x,\sigma],$ where $\sigma$ is the Frobenius automorphism of $\mathbb{F}_{q},$ and with $l,h$ are integers such that $150<l<h<$ 15. Let $\vec{a}(x):=cx^{h}+bx^{l}+a,$ the Hamming weight of $\vec{a}(x)$ is $3,$ then according to Theorem 6.4, the number of $(15,\sigma)$-equivalence classes is
$N=\frac{3^{3}}{\operatorname{lcm}(\frac{3}{\gcd([15]_{1},3)},\frac{3}{\gcd(2^{l}[15-l]_{1},3)},\frac{3}{\gcd(2^{h}[15-h]_{1},3)})}=\frac{3^{3}}{\operatorname{lcm}(1,\frac{3}{\gcd(15-l,3)},\frac{3}{\gcd(15-h,3)})}.$
  1. 1.
    If $l$ and $h$ are multiples of $3$ then the number if equivalence classes is $N=3^{3}$ and the equivalence relation has no influence in this case.
  2. 2.
    Else, i.e., $l$ or $h$ is not a multiple of $3$ then $N=\dfrac{3^{3}}{3}=9$. Let suppose that $h$ is a multiple and $l$ not, then as
    $\gcd\left(\frac{3}{\gcd(15,3)},\frac{3}{\gcd(15-l,3)},\frac{3}{\gcd(15-h,3)}\right)=\gcd(1,3,1)=1,$
    then by Corollary 6.5, each polycyclic codes associated with a polynomial $f$ of the form $f=x^{15}-cx^{h}-bx^{l}-a,$ is equivalent to a polycyclic code associated with one of the following polynomials:
    $f_{i,j}=x^{15}-\xi^{i}x^{h}-x^{l}-\xi^{j},\quad i,j\in\{0,1,2\}.$
    We again found some optimal (for given $n,k$ and $q$) polycyclic codes, which we present in Table 2.
    Polynomials $f$ Generator Polynomial Parameters
    $x^{15}+x^{3}+x^{2}+\xi$ $x^{3}+x^{2}+\xi^{2}$ $[15,12,3]_{4}$
    $x^{15}+\xi x^{3}+x^{2}+\xi$ $x^{5}+\xi^{2}x^{3}+x+\xi^{2}$ $[15,10,4]_{4}$
    $x^{15}+\xi x^{3}+x^{2}+\xi$ $x^{4}+x^{3}+\xi x^{2}+\xi x+\xi^{2}$ $[15,11,4]_{4}$
    $x^{15}+x^{6}+x^{2}+\xi$ $x^{10}+x^{8}+\xi^{2}x^{7}+\xi x^{6}+\xi x^{5}+\xi^{2}x^{4}+\xi^{2}x^{3}+\xi x^{2}+\xi x+1$ $[15,5,8]_{4}$
    $x^{15}+\xi x^{12}+x^{2}+1$ $x^{6}+\xi x^{5}+x^{4}+\xi x^{3}+\xi x^{2}+\xi x+\xi^{2}$ $[15,9,5]_{4}$
    $x^{15}+\xi^{2}x^{12}+x+\xi^{2}$ $x^{4}+x^{3}+\xi^{2}x^{2}+\xi^{2}x+\xi$ $[15,11,4]_{4}$
    Table 2: Optimal polycyclic codes with best known parameters of length $n=15$ over $\mathbb{F}_{4}$.
Conclusion
In this paper we investigated the equivalence between classes of polycyclic codes associated with certain polynomials over the finite field $\mathbb{F}_{q}$. We began with the specific case of polycyclic codes associated with trinomials of the form $x^{n}-a_{\ell}x^{\ell}-a_{0}$, which we refer to as $\ell$-trinomial codes. We introduced an equivalence relation called $n$-equivalence, which extends the notion of $n$-equivalence known for constacyclic codes. We derived a formula for the number of $n$-equivalence classes and provided conditions under which two $\ell$-trinomial code families are equivalent. We then focused on $p^{\ell}$-trinomial codes of length $p^{\ell+r}$, where $p$ is the characteristic of $\mathbb{F}_{q}$ and $r$ an integer, and established further results in this special case. In the end, we generalized our results to general polycyclic codes. Furthermore, using our results on equivalence to restrict our search space, we gave some examples of both trinomial and more general polycyclic codes with optimal or best known parameters.
In future work, we will consider more general isometries of the form $\varphi_{\alpha}(x)=\alpha x^{k}$ for a chosen integers $k$ (in this paper, we took $k=1$) to reduce the number of equivalence classes and to use them for a more refined classification of polycyclic codes. Moreover, we will study explicit factorizations of trinomial polynomials and their applications in the construction of trinomial codes.
Acknowledgments
The authors thank Adrien Pasquereau, Abdelghefar Chibloun, and Prof. Nuh Aydin for their many helpful discussions and suggestions to improve this paper.
Declarations
The first author’s research is supported by a Swiss Government Excellence Scholarship (ESKAS), no: 2024.0504.